Find the equation of the tangent plane to the conicoid x2 +y2 = kz at the point (k, k,2k), where k is a constant. Represent the plane geometrically. Now take different values of k, including both positive and negative, and see how the shape of the conicoid changes.


1
Expert's answer
2021-06-04T12:44:09-0400

We have an equation of conicoid


"F(x,y,z) = x^2 + y^2 - kz"


And the point


"M(k,k,2k)"


Let’s find a derivative


"F'_x(x,y,z) = 2x; F'_x(M) = 2k""F'_y(x,y,z) = 2y; F'_y(M) = 2k""F'_z(x,y,z) = -k; F'_z(M) = -k"

So equation of tangent plane is:


"F'_x(M)*(x-k) + F'_y(M)*(y-k) + F'_z(M)*(z-2k) = 0""2k*(x-k) + 2k*(y-k)+(-k)*(z-2k) = 2kx + 2ky - kz -2k^2 = 0"

Answer:

"2kx + 2ky - kz -2k^2 = 0"

if k > 0



if k < 0